OpenStudy (anonymous):

solve for x: 77=33^x

OpenStudy (anonymous):

\[log_{33}77\]

OpenStudy (anonymous):

how do i type that into my calculator

OpenStudy (anonymous):

Some calculator doesn't have base 33 logs. You can use this method instead: 77=33^x log77=log(33^x) log77=xlog33 x=log77/log33